OpenStudy (anonymous):

explain how the graph of y=2sin x can be made to coincide with the graph of y =2cosx

OpenStudy (anonymous):

shift it to the right \(\frac{\pi}{2}\) units

OpenStudy (anonymous):

oh sorry you want to make the graph of sine in to cosine shift it to the LEFT \(\frac{\pi}{2}\) units by writing \[\sin(x+\frac{\pi}{2})\]